You need to make an aqueous solution of 0.221 M chromium(III) acetate for an experiment in lab, using a 500 mL volumetric flask. How much solid chromium(III) acetate should you add?

________ grams
1
Expert's answer
2019-03-14T04:00:21-0400

The mass of the chromium (III) acetate equals m = C × V × Mr = 0.221 M × 0.5 L × 229 g/mol = 25.3045 g.

Answer: 25.3045 g.